# Wave Network: An Ultra-Small Language Model

### v3

@@ 3.1.1 Global Semantics Representation
From a signal processing perspective, signals are often represented in polar coordinates and processed in the frequency domain. A signal’s amplitude represents its strength, and its phase describes its position relative to another signal within a cycle [17]. Referring to signal processing, we treat token embeddings as discrete signals in the frequency domain. Given an input text with $n$ tokens $input\_text=[\mathbf{w}_{1},\mathbf{w}_{2},\dots,\mathbf{w}_{m},\dots,\mathbf{w}_{n}]$:
$&\textit{dimension 1}\;\;\;\;\textit{dimension k}\\[-5.0pt] &\;\;\;\;\;\;\;\downarrow\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\downarrow\\[-5.0pt] \mathbf{w}_{1}&=\left[w_{1,1},w_{1,2},\dots,w_{1,k},\dots,w_{1,768}\right]\\ \mathbf{w}_{2}&=\left[w_{2,1},w_{2,2},\dots,w_{2,k},\dots,w_{2,768}\right]\\[-5.0pt] &\makebox[45.0pt][c]{\vdots}\makebox[10.0pt][c]{\vdots}\makebox[60.0pt][c]{\vdots}\makebox[25.0pt][c]{\vdots}\\[-5.0pt] \mathbf{w}_{j}&=\left[w_{j,1},w_{j,2},\dots,w_{j,k},\dots,w_{j,768}\right]\\[-5.0pt] &\makebox[45.0pt][c]{\vdots}\makebox[10.0pt][c]{\vdots}\makebox[60.0pt][c]{\vdots}\makebox[25.0pt][c]{\vdots}\\[-5.0pt] \mathbf{w}_{m}&=\left[w_{m,1},w_{m,2},\dots,w_{m,k},\dots,w_{m,768}\right]\\[-5.0pt] &\makebox[45.0pt][c]{\vdots}\makebox[10.0pt][c]{\vdots}\makebox[60.0pt][c]{\vdots}\makebox[25.0pt][c]{\vdots}\\[-5.0pt] \mathbf{w}_{n}&=\left[w_{n,1},w_{n,2},\dots,w_{n,k},\dots,w_{n,768}\right]\\ &\;\;\;\;\textbf{Token Embedding Matrix}\end{aligned}\hskip 28.45274pt\begin{aligned} &\textit{dimension 1}\;\;\textit{dimension k}\\[-5.0pt] &\;\;\;\;\downarrow\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\downarrow\\[-5.0pt] &\left[G_{1,1},G_{1,2},\dots,G_{1,k},\dots,G_{1,768}\right]\\ &\left[G_{2,1},G_{2,2},\dots,G_{2,k},\dots,G_{2,768}\right]\\[-5.0pt] &\makebox[30.0pt][c]{\vdots}\makebox[20.0pt][c]{\vdots}\makebox[60.0pt][c]{\vdots}\makebox[25.0pt][c]{\vdots}\\[-5.0pt] &\left[G_{j,1},G_{j,2},\dots,G_{j,k},\dots,G_{j,768}\right]\\[-5.0pt] &\makebox[30.0pt][c]{\vdots}\makebox[20.0pt][c]{\vdots}\makebox[60.0pt][c]{\vdots}\makebox[25.0pt][c]{\vdots}\\[-5.0pt] &\left[G_{m,1},G_{m,2},\dots,G_{m,k},\dots,G_{m,768}\right]\\[-5.0pt] &\makebox[30.0pt][c]{\vdots}\makebox[20.0pt][c]{\vdots}\makebox[60.0pt][c]{\vdots}\makebox[25.0pt][c]{\vdots}\\[-5.0pt] &\left[G_{n,1},G_{n,2},\dots,G_{n,k},\dots,G_{n,768}\right]\\ &\;\;\;\;\;\;\textbf{Token Magnitude Matrix}\end{aligned}\vspace{-9pt}$ (1)
@@
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\mathbf{G}$                        $\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\text{G}_{1},\;\;\;\;\;\;\;\;\;\;\;\text{G}_{2},\;\;\;\dots\;,\;\;\;\;\;\text{G}_{k},\;\;\;\dots\;,\;\;\;\;\;\text{G}_{768}$ $\begin{array}[]{cccccccc}{\begin{array}[]{c}\mathbf{w}_{1}=\\ \\ \mathbf{w}_{2}=\\ \\[-5.0pt] \vdots\\[-5.0pt] \\ \mathbf{w}_{j}=\\ \\[-5.0pt] \vdots\\[-5.0pt] \\ \mathbf{w}_{m}=\\ \\[-5.0pt] \vdots\\[-5.0pt] \\ \mathbf{w}_{n}=\end{array}}&\sqrt{\begin{array}[]{c}w_{1,1}^{2}\\ \;+\\ w_{2,1}^{2}\\ \;+\\[-5.0pt] \vdots\\[-5.0pt] \;+\\ w_{j,1}^{2}\\ \;+\\[-5.0pt] \vdots\\[-5.0pt] \;+\\ w_{m,1}^{2}\\ \;+\\[-5.0pt] \vdots\\[-5.0pt] \;+\\ w_{n,1}^{2}\end{array}}&\begin{array}[]{c}\\ \\ \\ \\ \\ \\ \\ \\ \\ \\ \\ \\ \\ \\ \end{array}&\sqrt{\begin{array}[]{c}w_{1,2}^{2}\\ \;+\\ w_{2,2}^{2}\\ \;+\\[-5.0pt] \vdots\\[-5.0pt] \;+\\ w_{j,2}^{2}\\ \;+\\[-5.0pt] \vdots\\[-5.0pt] \;+\\ w_{m,2}^{2}\\ \;+\\[-5.0pt] \vdots\\[-5.0pt] \;+\\ w_{n,2}^{2}\end{array}}&\begin{array}[]{c}\dots\\ \\ \dots\\ \\[-5.0pt] \vdots\\[-5.0pt] \\ \dots\\ \\[-5.0pt] \vdots\\[-5.0pt] \\ \dots\\ \\[-5.0pt] \vdots\\[-5.0pt] \\ \dots\end{array}&\sqrt{\begin{array}[]{c}w_{1,k}^{2}\\ \;+\\ w_{2,k}^{2}\\ \;+\\[-5.0pt] \vdots\\[-5.0pt] \;+\\ w_{j,k}^{2}\\ \;+\\[-5.0pt] \vdots\\[-5.0pt] \;+\\ w_{m,k}^{2}\\ \;+\\[-5.0pt] \vdots\\[-5.0pt] \;+\\ w_{n,k}^{2}\end{array}}&\begin{array}[]{c}\dots\\ \\ \dots\\ \\[-5.0pt] \vdots\\[-5.0pt] \\ \dots\\ \\[-5.0pt] \vdots\\[-5.0pt] \\ \dots\\ \\[-5.0pt] \vdots\\[-5.0pt] \\ \dots\end{array}&\sqrt{\begin{array}[]{c}w_{1,768}^{2}\\ \;+\\ w_{2,768}^{2}\\ \;+\\[-5.0pt] \vdots\\[-5.0pt] \;+\\ w_{j,768}^{2}\\ \;+\\[-5.0pt] \vdots\\[-5.0pt] \;+\\ w_{m,768}^{2}\\ \;+\\[-5.0pt] \vdots\\[-5.0pt] \;+\\ w_{n,768}^{2}\end{array}}\end{array}$
@@
Next, we sum the magnitudes of all token embedding components along each dimension to define the global semantics vector $\mathbf{G}=[{G}_{1},{G}_{2},\dots,{G}_{k},\dots,{G}_{768}]$, where each global semantic element ${G_{k}}$ can be defined as $\text{G}_{k}=\left\|\mathbf{w}_{:,k}\right\|_{2}=\left\|[w_{1,k},w_{2,k},\dots,w_{j,k},\dots+w_{m,k},\dots,w_{n,k}]\right\|_{2}=\sqrt{{w_{1,k}^{2}+w_{2,k}^{2}+\dots+w_{j,k}^{2}+\dots+w_{m,k}^{2}+\dots+w_{n,k}^{2}}}$, as shown in Figure 2. Here, $w_{j,k}$ represents the $k$-th dimension of the $j$-th token embedding. This global semantic vector $\mathbf{G}$ represents the global semantics of the entire input text and will serve as the magnitude of the complex vector token representation of each token in polar coordinates.
For simplicity, we focus on input-level global semantics in this research. As shown in Figure 1, given a input text with $n$ token embeddings, the input-level global semantics vector can be defined as $\mathbf{input\_G}=\left[{input\_G}_{1},{input\_G}_{2},\dots,{input\_G}_{k},\dots,{input\_G}_{768}\right]$, where ${input\_G}_{k}=\sqrt{{w_{1,k}^{2}+w_{2,k}^{2}+\dots+\dots+w_{j,k}^{2}+\dots+w_{m,k}^{2}+\dots+w_{n,k}^{2}}}$.
@@ 3.1.2 Local Semantics Representation
$\mathbf{w}_{1}&=\left[w_{1,1},w_{1,2},\dots,w_{1,k},\dots,w_{1,768}\right]\\ \mathbf{w}_{2}&=\left[w_{2,1},w_{2,2},\dots,w_{2,k},\dots,w_{2,768}\right]\\[-5.0pt] &\makebox[45.0pt][c]{\vdots}\makebox[10.0pt][c]{\vdots}\makebox[60.0pt][c]{\vdots}\makebox[25.0pt][c]{\vdots}\\[-5.0pt] \mathbf{w}_{j}&=\left[w_{j,1},w_{j,2},\dots,w_{j,k},\dots,w_{j,768}\right]\\[-5.0pt] &\makebox[45.0pt][c]{\vdots}\makebox[10.0pt][c]{\vdots}\makebox[60.0pt][c]{\vdots}\makebox[25.0pt][c]{\vdots}\\[-5.0pt] \mathbf{w}_{m}&=\left[w_{m,1},w_{m,2},\dots,w_{m,k},\dots,w_{m,768}\right]\\[-5.0pt] &\makebox[45.0pt][c]{\vdots}\makebox[10.0pt][c]{\vdots}\makebox[60.0pt][c]{\vdots}\makebox[25.0pt][c]{\vdots}\\[-5.0pt] \mathbf{w}_{n}&=\left[w_{n,1},w_{n,2},\dots,w_{n,k},\dots,w_{n,768}\right]\\ &\;\;\;\;\textbf{Token Embedding Matrix}\end{aligned}\hskip 28.45274pt\begin{aligned} \boldsymbol{\alpha}_{1}&=\left[\alpha_{1,1},\alpha_{1,2},\dots,\alpha_{1,k},\dots,\alpha_{1,768}\right]\\ \boldsymbol{\alpha}_{2}&=\left[\alpha_{2,1},\alpha_{2,2},\dots,\alpha_{2,k},\dots,\alpha_{2,768}\right]\\[-5.0pt] &\makebox[45.0pt][c]{\vdots}\makebox[10.0pt][c]{\vdots}\makebox[60.0pt][c]{\vdots}\makebox[25.0pt][c]{\vdots}\\[-5.0pt] \boldsymbol{\alpha}_{j}&=\left[\alpha_{j,1},\alpha_{j,2},\dots,\alpha_{j,k},\dots,\alpha_{j,768}\right]\\[-5.0pt] &\makebox[45.0pt][c]{\vdots}\makebox[10.0pt][c]{\vdots}\makebox[60.0pt][c]{\vdots}\makebox[25.0pt][c]{\vdots}\\[-5.0pt] \boldsymbol{\alpha}_{m}&=\left[\alpha_{m,1},\alpha_{m,2},\dots,\alpha_{m,k},\dots,\alpha_{m,768}\right]\\[-5.0pt] &\makebox[45.0pt][c]{\vdots}\makebox[10.0pt][c]{\vdots}\makebox[60.0pt][c]{\vdots}\makebox[25.0pt][c]{\vdots}\\[-5.0pt] \boldsymbol{\alpha}_{n}&=\left[\alpha_{n,1},\alpha_{n,2},\dots,\alpha_{n,k},\dots,\alpha_{n,768}\right]\\ &\;\;\;\;\;\;\;\;\textbf{Token Phase Matrix}\end{aligned}\vspace{-11pt}$ (2)
@@ 3.2.1 Wave Interference
From physical perspective, wave interference is a phenomenon where two coherent waves combined by adding their intensities or displacements, considering their phase difference. In the context of multiple levels of global semantics, as discussed in the Subsection 3.1.2, we consider two variants of complex vector token representations for token $\boldsymbol{w_{j}}$ as: $\mathbf{input\_Z_{j}}=\mathbf{input\_G}\cdot e^{i\cdot\boldsymbol{input\_\alpha_{j}}}$ and $\mathbf{input\_Z_{j}^{\prime}}=\mathbf{input\_G^{\prime}}\cdot e^{i\cdot\boldsymbol{input\_\alpha_{j}^{\prime}}}$. We use complex vectors addition to simulate wave interference [22] and obtain the combined complex vector token representation $\mathbf{interference\_Z_{j}}$ for token $\boldsymbol{w_{j}}$ as follows:
@@
$\begin{split}&{Interference\_Z_{j,k}}=\sqrt{{w_{1,k}^{2}+w_{2,k}^{2}+\dots+w_{j,k}^{2}+\dots+w_{m,k}^{2}+\dots+w_{n,k}^{2}}}\\ &\cdot\frac{w_{j,k}}{\sqrt{{w_{j,k}^{2}+w_{2,k}^{2}+\dots+\dots+w_{j,k}^{2}+\dots+w_{m,k}^{2}+\dots+w_{n,k}^{2}}}}\\ &+\sqrt{{w_{1,k}^{\prime 2}+w_{2,k}^{\prime 2}+\dots+w_{j,k}^{\prime 2}+\dots+w_{m,k}^{\prime 2}+\dots+w_{n,k}^{\prime 2}}}\cdot\frac{w_{j,k}^{\prime}}{\sqrt{{w_{j,k}^{\prime 2}+w_{2,k}^{\prime 2}+\dots+\dots+w_{j,k}^{\prime 2}+\dots+w_{m,k}^{\prime 2}+\dots+w_{n,k}^{\prime 2}}}}\\ &+i\cdot(\sqrt{{w_{1,k}^{2}+w_{2,k}^{2}+\dots+w_{j,k}^{2}+\dots+w_{m,k}^{2}+\dots+w_{n,k}^{2}}})\\ &\cdot(\sqrt{1-(\frac{w_{j,k}}{\sqrt{{w_{1,k}^{2}+w_{2,k}^{2}+\dots+\dots+w_{j,k}^{2}+\dots+w_{m,k}^{2}+\dots+w_{n,k}^{2}}}})^{2}})\\ &+i\cdot(\sqrt{{w_{1,k}^{\prime 2}+w_{2,k}^{\prime 2}+\dots+w_{j,k}^{\prime 2}+\dots+w_{m,k}^{\prime 2}+\dots+w_{n,k}^{\prime 2}}})\\ &\cdot(\sqrt{1-(\frac{w_{j,k}^{\prime}}{\sqrt{{w_{1,k}^{\prime 2}+w_{2,k}^{\prime 2}+\dots+\dots+w_{j,k}^{\prime 2}+\dots+w_{m,k}^{\prime 2}+\dots+w_{n,k}^{\prime 2}}}})^{2}})\\ &=w_{j,k}+w^{\prime}_{j,k}+i\cdot(\sqrt{{w_{1,k}^{2}+w_{2,k}^{2}+\dots+w_{j-1,k}^{2}+w_{j+1,k}^{2}+\dots+w_{m,k}^{2}+\dots+w_{n,k}^{2}}}\\ &+\sqrt{{w_{1,k}^{\prime 2}+w_{2,k}^{\prime 2}+\dots+w_{j-1,k}^{\prime 2}+w_{j+1,k}^{\prime 2}+\dots+w_{m,k}^{\prime 2}+\dots+w_{n,k}^{\prime 2}}})\end{split}$ (4)

### v4

@@ 3.1.1 Global Semantics Representation
From a signal processing perspective, signals are often represented in polar coordinates and processed in the frequency domain. A signal’s amplitude represents its strength, and its phase describes its position relative to another signal within a cycle [17]. Referring to signal processing, we treat token embeddings as discrete signals in the frequency domain. Given an input text with $n$ tokens $input\_text=[\mathbf{w}_{1},\mathbf{w}_{2},\dots,\mathbf{w}_{j},\dots,\mathbf{w}_{n}]$:
$&\textit{dimension 1}\;\;\;\;\textit{dimension k}\\[-5.0pt] &\;\;\;\;\;\;\;\downarrow\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\downarrow\\[-5.0pt] \mathbf{w}_{1}&=\left[w_{1,1},w_{1,2},\dots,w_{1,k},\dots,w_{1,768}\right]\\ \mathbf{w}_{2}&=\left[w_{2,1},w_{2,2},\dots,w_{2,k},\dots,w_{2,768}\right]\\[-5.0pt] &\makebox[45.0pt][c]{\vdots}\makebox[10.0pt][c]{\vdots}\makebox[60.0pt][c]{\vdots}\makebox[25.0pt][c]{\vdots}\\[-5.0pt] \mathbf{w}_{j}&=\left[w_{j,1},w_{j,2},\dots,w_{j,k},\dots,w_{j,768}\right]\\[-5.0pt] &\makebox[45.0pt][c]{\vdots}\makebox[10.0pt][c]{\vdots}\makebox[60.0pt][c]{\vdots}\makebox[25.0pt][c]{\vdots}\\[-5.0pt] \mathbf{w}_{n}&=\left[w_{n,1},w_{n,2},\dots,w_{n,k},\dots,w_{n,768}\right]\\ &\;\;\;\;\textbf{Token Embedding Matrix}\end{aligned}\hskip 28.45274pt\begin{aligned} &\textit{dimension 1}\;\;\textit{dimension k}\\[-5.0pt] &\;\;\;\;\downarrow\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\downarrow\\[-5.0pt] &\left[G_{1,1},G_{1,2},\dots,G_{1,k},\dots,G_{1,768}\right]\\ &\left[G_{2,1},G_{2,2},\dots,G_{2,k},\dots,G_{2,768}\right]\\[-5.0pt] &\makebox[30.0pt][c]{\vdots}\makebox[20.0pt][c]{\vdots}\makebox[60.0pt][c]{\vdots}\makebox[25.0pt][c]{\vdots}\\[-5.0pt] &\left[G_{j,1},G_{j,2},\dots,G_{j,k},\dots,G_{j,768}\right]\\[-5.0pt] &\makebox[30.0pt][c]{\vdots}\makebox[20.0pt][c]{\vdots}\makebox[60.0pt][c]{\vdots}\makebox[25.0pt][c]{\vdots}\\[-5.0pt] &\left[G_{n,1},G_{n,2},\dots,G_{n,k},\dots,G_{n,768}\right]\\ &\;\;\;\;\;\;\textbf{Token Magnitude Matrix}\end{aligned}\vspace{-9pt}$ (1)
@@
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\mathbf{G}$                        $\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\text{G}_{1},\;\;\;\;\;\;\;\;\;\;\;\text{G}_{2},\;\;\;\dots\;,\;\;\;\;\;\text{G}_{k},\;\;\;\dots\;,\;\;\;\;\;\text{G}_{768}$ $\begin{array}[]{cccccccc}{\begin{array}[]{c}\mathbf{w}_{1}=\\ \\ \mathbf{w}_{2}=\\ \\[-5.0pt] \vdots\\[-5.0pt] \\ \mathbf{w}_{j}=\\ \\[-5.0pt] \vdots\\[-5.0pt] \\ \mathbf{w}_{n}=\end{array}}&\sqrt{\begin{array}[]{c}w_{1,1}^{2}\\ \;+\\ w_{2,1}^{2}\\ \;+\\[-5.0pt] \vdots\\[-5.0pt] \;+\\ w_{j,1}^{2}\\ \;+\\[-5.0pt] \vdots\\[-5.0pt] \;+\\ w_{n,1}^{2}\end{array}}&\begin{array}[]{c}\\ \\ \\ \\ \\ \\ \\ \\ \\ \\ \\ \end{array}&\sqrt{\begin{array}[]{c}w_{1,2}^{2}\\ \;+\\ w_{2,2}^{2}\\ \;+\\[-5.0pt] \vdots\\[-5.0pt] \;+\\ w_{j,2}^{2}\\ \;+\\[-5.0pt] \vdots\\[-5.0pt] \;+\\ w_{n,2}^{2}\end{array}}&\begin{array}[]{c}\dots\\ \\ \dots\\ \\[-5.0pt] \vdots\\[-5.0pt] \\ \dots\\ \\[-5.0pt] \vdots\\[-5.0pt] \\ \dots\end{array}&\sqrt{\begin{array}[]{c}w_{1,k}^{2}\\ \;+\\ w_{2,k}^{2}\\ \;+\\[-5.0pt] \vdots\\[-5.0pt] \;+\\ w_{j,k}^{2}\\ \;+\\[-5.0pt] \vdots\\[-5.0pt] \;+\\ w_{n,k}^{2}\end{array}}&\begin{array}[]{c}\dots\\ \\ \dots\\ \\[-5.0pt] \vdots\\[-5.0pt] \\ \dots\\ \\[-5.0pt] \vdots\\[-5.0pt] \\ \dots\end{array}&\sqrt{\begin{array}[]{c}w_{1,768}^{2}\\ \;+\\ w_{2,768}^{2}\\ \;+\\[-5.0pt] \vdots\\[-5.0pt] \;+\\ w_{j,768}^{2}\\ \;+\\[-5.0pt] \vdots\\[-5.0pt] \;+\\ w_{n,768}^{2}\end{array}}\end{array}$
@@
Next, we sum the magnitudes of all token embedding components along each dimension to define the global semantics vector $\mathbf{G}=[{G}_{1},{G}_{2},\dots,{G}_{k},\dots,{G}_{768}]$, where each global semantic element ${G_{k}}$ can be defined as $\text{G}_{k}=\left\|\mathbf{w}_{:,k}\right\|_{2}=\left\|[w_{1,k},w_{2,k},\dots,w_{j,k},\dots,w_{n,k}]\right\|_{2}=\sqrt{{w_{1,k}^{2}+w_{2,k}^{2}+\dots+w_{j,k}^{2}+\dots+w_{n,k}^{2}}}$, as shown in Figure 2. Here, $w_{j,k}$ represents the $k$-th dimension of the $j$-th token embedding. This global semantic vector $\mathbf{G}$ represents the global semantics of the entire input text and will serve as the magnitude of the complex vector token representation of each token in polar coordinates.
For simplicity, we focus on input-level global semantics in this research. As shown in Figure 1, given a input text with $n$ token embeddings, the input-level global semantics vector can be defined as $\mathbf{input\_G}=\left[{input\_G}_{1},{input\_G}_{2},\dots,{input\_G}_{k},\dots,{input\_G}_{768}\right]$, where ${input\_G}_{k}=\sqrt{{w_{1,k}^{2}+w_{2,k}^{2}+\dots+w_{j,k}^{2}+\dots+w_{n,k}^{2}}}$.
@@ 3.1.2 Local Semantics Representation
$\mathbf{w}_{1}&=\left[w_{1,1},w_{1,2},\dots,w_{1,k},\dots,w_{1,768}\right]\\ \mathbf{w}_{2}&=\left[w_{2,1},w_{2,2},\dots,w_{2,k},\dots,w_{2,768}\right]\\[-5.0pt] &\makebox[45.0pt][c]{\vdots}\makebox[10.0pt][c]{\vdots}\makebox[60.0pt][c]{\vdots}\makebox[25.0pt][c]{\vdots}\\[-5.0pt] \mathbf{w}_{j}&=\left[w_{j,1},w_{j,2},\dots,w_{j,k},\dots,w_{j,768}\right]\\[-5.0pt] &\makebox[45.0pt][c]{\vdots}\makebox[10.0pt][c]{\vdots}\makebox[60.0pt][c]{\vdots}\makebox[25.0pt][c]{\vdots}\\[-5.0pt] \mathbf{w}_{n}&=\left[w_{n,1},w_{n,2},\dots,w_{n,k},\dots,w_{n,768}\right]\\ &\;\;\;\;\textbf{Token Embedding Matrix}\end{aligned}\hskip 28.45274pt\begin{aligned} \boldsymbol{\alpha}_{1}&=\left[\alpha_{1,1},\alpha_{1,2},\dots,\alpha_{1,k},\dots,\alpha_{1,768}\right]\\ \boldsymbol{\alpha}_{2}&=\left[\alpha_{2,1},\alpha_{2,2},\dots,\alpha_{2,k},\dots,\alpha_{2,768}\right]\\[-5.0pt] &\makebox[45.0pt][c]{\vdots}\makebox[10.0pt][c]{\vdots}\makebox[60.0pt][c]{\vdots}\makebox[25.0pt][c]{\vdots}\\[-5.0pt] \boldsymbol{\alpha}_{j}&=\left[\alpha_{j,1},\alpha_{j,2},\dots,\alpha_{j,k},\dots,\alpha_{j,768}\right]\\[-5.0pt] &\makebox[45.0pt][c]{\vdots}\makebox[10.0pt][c]{\vdots}\makebox[60.0pt][c]{\vdots}\makebox[25.0pt][c]{\vdots}\\[-5.0pt] \boldsymbol{\alpha}_{n}&=\left[\alpha_{n,1},\alpha_{n,2},\dots,\alpha_{n,k},\dots,\alpha_{n,768}\right]\\ &\;\;\;\;\;\;\;\;\textbf{Token Phase Matrix}\end{aligned}\vspace{-11pt}$ (2)
@@ 3.2.1 Wave Interference
From physical perspective, wave interference is a phenomenon where two coherent waves combined by adding their intensities or displacements, considering their phase difference. In the context of generating complex vector token representations from input-level global semantics, as discussed in the Subsection 3.1.1, we consider two variants of complex vector token representations for token $\boldsymbol{w_{j}}$ as: $\mathbf{input\_Z_{j}}=\mathbf{input\_G}\cdot e^{i\cdot\boldsymbol{input\_\alpha_{j}}}$ and $\mathbf{input\_Z_{j}^{\prime}}=\mathbf{input\_G^{\prime}}\cdot e^{i\cdot\boldsymbol{input\_\alpha_{j}^{\prime}}}$. We use complex vectors addition to simulate wave interference [22] and obtain the combined complex vector token representation $\mathbf{interference\_Z_{j}}$ for token $\boldsymbol{w_{j}}$ as follows:
@@
$\begin{split}&{Interference\_Z_{j,k}}=\sqrt{{w_{1,k}^{2}+w_{2,k}^{2}+\dots+w_{j,k}^{2}+\dots+w_{n,k}^{2}}}\cdot\frac{w_{j,k}}{\sqrt{{w_{j,k}^{2}+w_{2,k}^{2}+\dots+\dots+w_{j,k}^{2}+\dots+w_{n,k}^{2}}}}\\ &+\sqrt{{w_{1,k}^{\prime 2}+w_{2,k}^{\prime 2}+\dots+w_{j,k}^{\prime 2}+\dots+w_{n,k}^{\prime 2}}}\cdot\frac{w_{j,k}^{\prime}}{\sqrt{{w_{j,k}^{\prime 2}+w_{2,k}^{\prime 2}+\dots+\dots+w_{j,k}^{\prime 2}+\dots+w_{n,k}^{\prime 2}}}}\\ &+i\cdot(\sqrt{{w_{1,k}^{2}+w_{2,k}^{2}+\dots+w_{j,k}^{2}+\dots+w_{n,k}^{2}}})\cdot(\sqrt{1-(\frac{w_{j,k}}{\sqrt{{w_{1,k}^{2}+w_{2,k}^{2}+\dots+w_{j,k}^{2}+\dots+w_{n,k}^{2}}}})^{2}})\\ &+i\cdot(\sqrt{{w_{1,k}^{\prime 2}+w_{2,k}^{\prime 2}+\dots+w_{j,k}^{\prime 2}+\dots+w_{n,k}^{\prime 2}}})\cdot(\sqrt{1-(\frac{w_{j,k}^{\prime}}{\sqrt{{w_{1,k}^{\prime 2}+w_{2,k}^{\prime 2}+\dots+w_{j,k}^{\prime 2}+\dots+w_{n,k}^{\prime 2}}}})^{2}})\\ &=w_{j,k}+w^{\prime}_{j,k}+i\cdot(\sqrt{{w_{1,k}^{2}+w_{2,k}^{2}+\dots+w_{j-1,k}^{2}+w_{j+1,k}^{2}+\dots+w_{n,k}^{2}}}\\ &+\sqrt{{w_{1,k}^{\prime 2}+w_{2,k}^{\prime 2}+\dots+w_{j-1,k}^{\prime 2}+w_{j+1,k}^{\prime 2}+\dots+w_{n,k}^{\prime 2}}})\end{split}$ (4)